The reporting period commenced with a relatively low activity rate. This increased by the middle of the period and then decreased again towards the end. Seismicity has been dominated by rockfall signals, although hybrid earthquakes have been seen at elevated levels, continuing from the end of the last reporting period. Dome growth has continued and became more concentrated to the southern flank of the dome as the week progressed.
Viewing conditions were generally poor during most of the week because of either low cloud cover on the volcano and/or rain. However, there were intermittent periods during which the dome was visible. Aerial views in the early part of the week showed that material had continued to pile up against the Gages and Farrell's walls. However, only a small amount of material had spilled over the Gages Wall while the moat area behind Farrell's Wall was still about 75 ft deep. A small spine which had grown on the northeast flank of the dome collapsed around 12/13 June. Two small debris flows from both north and south of Castle Peak were observed by the Airport Control Tower in the early morning of 14 June and the associated ash plumes were visible from the Observatory. Near-continuous ash production was noted between about 18:00 and 19:00 on 14 June at a time of high rockfall activity and also heavy rain. In the early afternoon of 15 June, a small pyroclastic flow was seen occurring in the south fork of the Tar River to the south of Castle Peak and a later helicopter flight revealed that this flow travelled about 1,000 ft (300 m) from the dome. Other deposits from small pyroclastic flows were noted in the Upper Tar River valley, both from the eastern and northeastern flanks of the dome.
A series of small pyroclastic flows were seen from Bramble Airport starting at 09:52 on 16 June, travelling down the valley on the southern side of Castle Peak. The first travelled about 2,500 ft from the dome, setting fire to a few trees. This was larger than any previous flows in this area. The second flow, which occurred at 09:57, did not travel as far as the first. There then followed a series of smaller flows. The larger flows generated a moderately thick ash cloud which drifted westwards on the wind, depositing ash on Plymouth and the surrounding areas. A helicopter inspection showed that these flows were not as hot as previous ones since many trees were covered in ash, but not burnt. There were further small pyroclastic flows seen in the area south of Castle Peak at 10:21, 11:40 and 13:07.
Observations from the helicopter in the later part of the week revealed that rockfall activity continued to some extent on the northeastern, northern, and southeastern flanks of the dome, but at slightly higher rates in the south and southwestern sectors. Material in the moat beneath Farrell's Wall had piled up to a height of about 50 ft from the top. However, very little new material was found to have fallen into the Upper Fort Ghaut area over the Gages Wall.
Seismicity throughout the week continued to be dominated by signals attributed to rockfalls from the growing dome (Table 1). In particular, there was an increase in the numbers of rockfall signals from about 14:00 on 14 June and some of the signals from events on this day were probably due to small pyroclastic flows. The elevated level of rockfall activity continued through 16 June and rockfall signals were almost continuous between 09:48 and 10:05 on 16 June.
Table 1. Earthquake types 13 June - 19 June 1996. These counts are taken from analysis of events triggered by the seismic network and do not necessarily coincide with numbers recorded in daily reports, which are taken from helicorder records.
Date VT LP Hybrid Rockfalls Tremor Amount 13/06 1 1 39 36 Low 14/06 1 19 61 114 Low 15/06 0 6 77 157 N/A 16/06 0 22 7 151 N/A 17/06 1 15 4 77 Low 18/06 0 8 1 83 Low to intermediate 19/06 0 5 7 82 Intermediate
MGAT (Gages) and MCPZ (Chances Peak) began recording a few small repetitive hybrid events very early in the morning of 13 June and these slowly increased in number and were occurring at a rate of about 1 every 2 minutes by 16:00 on 13 June. By the afternoon (c. 14:00) of 14 June, the rate of occurrence of these events was about one every five minutes. It increased during the night to about one per minute, then decreased again to about one per four minutes in the afternoon of 15 June. These events continued at a declining rate but with slightly increased amplitude, and by the afternoon of 16 June, they were occurring at a rate of only two or three per hour. There was a continuing decline in the number and size of repetitive hybrid events and these had completely died out by 18 June.
Two volcano-tectonic earthquakes were recorded and located during this reporting period; one on 13 June beneath the crater at a depth of 1 km and the other on 18 June at a depth of 6 km, located approximately 12 km NE of Windy Hill.
Several episodes of intermittent low- to moderate-amplitude broadband tremor were recorded on the closest stations to the volcano (MGAT and MCPZ) and these were most probably attributable to increased steam emission from the dome.
The eastern, northern and western EDM triangles were reoccupied at least once during the weak. The lines to Castle Peak , from both Long Ground and Whites, showed shortening during the period of the order of about 5 mm per day, consistent with the results of trends during the past 3 weeks. The changes in line length to the Chance's Peak steps were consistent with past results, with movements of less than 1 mm per day recorded. No changes in line lengths were observed for the northern triangle.
The MVO GPS programme was restarted after replacement of faulty equipment and measurements were done on the western, northern and eastern flanks of the volcano during the week. The results seem to show no changes above the error of the method, although a long term baseline is still required before more detailed modelling can be undertaken.
No tilt events were recorded by the Long Ground tiltmeter.
COSPEC measurements were made three times during the week and the results obtained were generally low, with SO2 flux ranging from 125 to 170 tonnes per day. The values are consistent with the generally low levels which have been recorded since COSPEC measurements restarted in late April 1996.
Collection of ash and water samples and monitoring of SO2 continues in areas around the volcano; these tests give results through which the environmental and health impacts of the volcano can be monitored.
Jean-Pierre Viode, Observatoire de la Montagne Pelee, Martinique.
Paul Cole, University of Luton, UK.
Christian Antenor-Habazac, Observatoire de la Soufriere, Guadeloupe.